# Controlling Space and Time with Diffusion Models

### v1

@@ Abstract
We present 4DiM, a cascaded diffusion model for 4D novel view synthesis (NVS), conditioned on one or more images of a general scene, and a set of camera poses and timestamps. To overcome challenges due to limited availability of 4D training data, we advocate joint training on 3D (with camera pose), 4D (pose+time) and video (time but no pose) data and propose a new architecture that enables the same. We further advocate the calibration of SfM posed data using monocular metric depth estimators for metric scale camera control. For model evaluation, we introduce new metrics to enrich and overcome shortcomings of current evaluation schemes, demonstrating state-of-the-art results in both fidelity and pose control compared to existing diffusion models for 3D NVS, while at the same time adding the ability to handle temporal dynamics. 4DiM is also used for improved panorama stitching, pose-conditioned video to video translation, and several other tasks.
@@
| Input 1 | Input 2 | Extrapolate time forward (waves approaching) |
| Extrapolate time backward (waves receding) | ||
@@ 1 Introduction
Novel view synthesis (NVS) and 3D generative models have emerged as a new frontier in generative models, enabling the synthesis of novel views of 3D objects and scenes with control over camera pose. These models compliment text-to-image/video models, with potential uses in the generation of 3D assets, augmented reality, synthetic data for model training (e.g., in robotics), and view interpolation. Notable recent examples using diffusion models include 3DiM (Watson et al., 2022) and Zero-1-to-3 (Liu et al., 2023b). They do not do generate explicit 3D scenes, rather, they leverage implicit 3D knowledge for conditional view generation, as a form of image+pose conditioned image generation.
To date, these models have focused mainly on objects with blank backgrounds, rather than generic natural scenes, and on limited camera poses, like those over the surface of the view-sphere fixated on objects centered at the origin. While recent work (Wang et al., 2023a; Yu et al., 2023a) has attempted to train models for generic scenes and 3D viewpoints, zero-shot performance on out of distribution scenes remains fraught. The biggest challenge is that 3D scene data is scarce. Much of the existing data also relies on COLMAP (Schönberger and Frahm, 2016; Schönberger et al., 2016) for the estimation of camera pose, yielding noisy poses with unknown scale. This is additionally problematic for effective control in the single-image to 3D task because models must sample from the distribution of plausible scales at inference time.
@@
The goal of this work is to extend NVS diffusion models in three respects: 1) from objects to scenes; 2) to free-form camera poses specified in meaningful physical units; and 3) to simultaneously allow spatial and temporal control via camera pose and timestamp conditioning. To this end, we introduce 4DiM, a diffusion model for NVS conditioned on (one or more) images of arbitrary scenes, camera pose, and time. 4DiM is trained on a mixture of data sources, including posed 3D images/video and unposed video, of both indoor and outdoor scenes. We make this possible through various technical innovations that enable training from missing data (e.g., images without pose or time annotations) and sampling with separate guidance weights on different conditioning images, poses or times. We also introduce calibrated versions of video datasets posed via COLMAP. Calibrated data makes it easier to learn metric regularities in the world, like the typical sizes of everyday objects and spatial relations, and it enables one to specify camera poses in meaningful physical units. As such, 4DiM generates 3D consistent, multi-view images or video of dynamic scenes. Beyond sample generation 4DiM can be used to unlock myriad applications, like video-to-video translation, improved panoramic stitching, or training explicit 3D models with score-distillation sampling (Poole et al., 2022).
To summarize, our main contributions include:
@@
4DiM, a pixel-based diffusion model for novel view synthesis conditioned on one or more images of arbitrary scenes, camera pose, and time. 4DiM comprises a base model that generates 32 images at , and a multi-view super resolution model that up-samples to ;
@@
An effective data mixture for 4D models comprising posed and unposed video of indoor and outdoor scenes, enabling zero-shot application with fine-grained pose control and dynamics;
Novel architectural elements to facilitate time and pose conditioning as well as training with incomplete data, and multi-guidance to simultaneously guide on camera pose and timestamps;
@@
A Calibrated version of RealEstate10K to improve model fidelity and enable metric pose control;
@@
Extensive evaluation and comparisons to prior work, including novel SfM-based metrics, namely, SfM distances and keypoint distance to better quantify pose alignment and dynamics.
@@ 2 Related work
We next provide a brief overview of prior work on camera control and 3D NVS with diffusion models. We also briefly discuss some recent work on 3D extraction using geometry-aware methods, which has emerged as an area of key interest to the research community. Here, the use of geometry-free pose-conditional diffusion models (like 4DiM) has been the key building block.
@@
Subsequent work proposed attention mechanisms leveraging epipolar geometry to improve the 3D consistency of image-to-image diffusion models for NVS (Tseng et al., 2023), and more recently, finetuning text-to-video models with temporal attention layers or attention layers limited to overlapping regions (He et al., 2024; Wang et al., 2023a) to model the joint distribution of generated views and their camera extrinsics. These models however still suffer from several issues: they have difficulty with static scenes due to persistent dynamics from the underlying video models; they still suffer from 3D inconsistencies and low fidelity; and they exhibit poor generalization to out-of-distribution image inputs. Alternatives have been proposed to improve fidelity in multi-view diffusion models of 3D scenes, albeit sacrificing the ability to model free-form camera pose; e.g., see MVDiffusion and follow-ups (Tang et al., 2023, 2024) where the data distribution is limited in its pose trajectories.
A concurrent trajectory of closely related work on 3D extraction has recently emerged following DreamFusion (Poole et al., 2022), where instead of directly training diffusion models for NVS, new techniques for sampling views parametrized as a NeRF with volume rendering are proposed such as Score Distillation Sampling (SDS) and Variational Score Distillation (Wang et al., 2024) (VSD). Here, a pre-existing diffusion model acts as a prior that drives the generation process. This enables, for example, text-to-3D using a text-to-image diffusion model. As demonstrated by MVDream (Shi et al., 2023), ReconFusion (Wu et al., 2023) and CAT3D (Gao* et al., 2024), using a pose-conditional diffusion model offers the unique advantage that the diffusion model can produce samples at the exact viewpoint specified for volume rendering during score distillation or NeRF postprocessing, which results in dramatically improved sample quality. All of these works establish further interest on improving pose-conditional diffusion models, which is the focus of our work, thereby enabling improvements in 3D extraction methods that build upon such models.
@@
4DiM uses a continuous-time diffusion model to learn the joint distribution over multiple views:
where are generated images, are conditioning images, are relative camera poses (extrinsics and intrinsics) and are scalar, relative timestamps.1114DiM does not require sequential temporal ordering as in video models as the architecture is permutation-equivariant over frames. All images (conditioning and generated) are processed by the diffusion model. Following Ho et al. (2020), we choose our loss function to be the error between the predicted and actual noise ( in their paper), though we use the L1 rather than L2 norm, because, like prior work (Saharia et al., 2022a; Saxena et al., 2023), we found that it improves sample quality. Our models use the “-parametrization” (Salimans and Ho, 2022), which helps stabilize training, and we adopt the noise schedules proposed by Kingma et al. (2021). Our current models process or 32 images at resolution ( is the number of conditioning plus generated frames). To this end, we decompose the task into two models similar to prior work (Ho et al., 2022a, b); we first generate images at , and then up-sample using a super-resolution model trained with noise conditioning augmentation (Saharia et al., 2022c). We also finetuned the 32-frame model to condition on 2 and 8 frames to enable more comparisons and example applications of 4DiM models.
@@
One particularly rich dataset for training 3D models is RealEstate10K (Zhou et al., 2018) (RE10K). It comprises 10,000 video segments of static scenes for which SfM (Schönberger and Frahm, 2016) has been used to infer per-frame camera pose, but only up to an unknown length scale. The lack of metric scale makes training more difficult because metric regularities of the world are lost, and it becomes more difficult for users to specify the target camera poses or camera motions in any intuitive units. This is especially problematic when conditioning on a single image, where scale itself otherwise becomes ambiguous. We therefore created a calibrated version of RealEstate10K by regressing the unknown metric scale from a monocular depth estimation model (Saxena et al., 2023). (For details, see Supplementary Material A.) The resulting dataset, cRE10K, has a major impact on model performance (see Section 4 below). We follow the same procedure to calibrate the LLFF dataset (Mildenhall et al., 2019) for evaluation purposes.
@@
Relatively little training data has both time and camera pose annotations; most 3D data represent static scenes, while video data rarely include camera pose. Finding a way to effectively condition on both camera pose and time in a way that allows for incomplete training data is essential. We thus propose to chain “Masked FiLM” layers (Dumoulin et al., 2018) for (positional encodings of) diffusion noise levels, per-pixel ray origins and directions, and video timestamps. When any of these conditioning signals is missing (due to incomplete training data or random dropout for classifier-free guidance (Ho and Salimans, 2022)), the FiLM layers are designed to reduce to the identity function, rather than simply setting missing values to zero. This avoids the network from confusing a timestamp of zero with dropped or missing data. In practice, we replace the FiLM shift with zeros and scale with ones. We illustrate the overall choices in Fig. 2. (For details see Supplementary Material B.)
@@
Steering 4DiM with the correct sampling hyperparameters is essential, especially the guidance weights for classifier-free guidance (Ho and Salimans, 2022) (CFG). In its usual formulation, CFG treats all conditioning variables as “one big variable”. In practice, placing a different weight on each variable is important; e.g., text requires high guidance weights, but high guidance weights on images can quickly lead to unwanted artifacts. We thus propose multi-guidance, where we generalize CFG to do exactly this, without making independence assumptions between conditioning variables. We start from a classifier-guided formulation (Dhariwal and Nichol, 2021), where we wish to sample with conditioning signals, , using the score
@@
If we train our model such that it drops out only , or and , … , or all of , we can sample with different guidance weights on each conditioning variable . 4DiM is trained to drop out conditioning signals with probability 0.1, and it drops either , and , or all of , and . This is a natural choice, since poses and timestamps without corresponding conditioning images do not convey useful information. For best results, 4DiM uses a guidance weight 2.0 on conditioning images, a weight of 4.0 on camera poses, and 1.0 for timestamps.
@@ 4 Evaluation
Evaluating 4D generative models is challenging. Typically, methods for NVS are evaluated using generation quality. In addition, for pose-conditioned generation it is desirable to find metrics that capture 3D consistency (rendered views should be grounded in a consistent 3D scene) and pose alignment (the camera should move as expected). Furthermore, evaluating temporal conditioning requires some measure that captures the motion of dynamic content. We leverage several existing metrics and propose new ones to cover all these aspects. Below we cover each one in detail.
@@
FID (Heusel et al., 2017) is a key metric for image quality, but if used in isolation it can be uninformative and a poor objective for hyper-parameter selection. For example, Saharia et al. (2022b) found that FID scores for text-to-image models are optimal with low CFG weights (Ho and Salimans, 2022), but text-image alignment suffers under his setting. In what follows we report FID as well as the improved Frechet distance with DINOv2 features (FDD) (Oquab et al., 2023), which appears to correlate with human judgements better than InceptionV3 features (Szegedy et al., 2016). For video quality, we also report FVD scores (Unterthiner et al., 2018).
@@
The work of Yu et al. (2023a) introduced the thresholded symmetric epipolar distance (TSED) as a more computationally efficient alternative to training separate NeRF models (Mildenhall et al., 2021) on every sample to evaluate 3D consistency (Watson et al., 2022). First, SIFT is used to obtain keypoints between a pair of views. For each keypoint in the first image, we can compute the epipolar line in the second image and its minimum distance to the corresponding keypoint. This is repeated for each keypoint in the second image to obtain a SED score. A hyperparameter threshold is selected, and the percentage of image pairs whose median SED is below the threshold is reported. Below, we report TSED with a threshold of 2.0 and include the TSED score of ground truth data, as poses in the data are noisy and do not achieve a perfect score.
Pose alignment (new metric: SfM distances)
We propose a set of metrics which we name SfM distances. Here, we run COLMAP pose estimation on the generated views, and compare the camera extrinsics predicted by COLMAP to the target poses. To get best results from COLMAP, we also feed it the input views. We report the relative error in camera positions (SfMDpos) and the angular deviation of camera rotations in radians (SfMDrot). Like we do for TSED scores, we also report the SfM distances for the ground truth data. Due to the inherent scale and rotational ambiguity of COLMAP, additional care must be taken to align the estimated poses to the original ones before comparing their differences as described above. Please see Supplementary Material D for more detail.
@@
One common failure mode that we observed early in our work is a tendency for certain models to copy input frames instead of generating temporal dynamics, and prior work has noted that good FVD scores can still be obtained with static video (Ge et al., 2024). We therefore propose a new metric called keypoint distance (KD), where we compute the average motion of SIFT keypoints(Lowe, 2004) across image pairs with matches. We report results on generated and reference views to assess whether generated images have a similar motion distribution.

### v2

@@ Abstract
We present 4DiM, a cascaded diffusion model for 4D novel view synthesis (NVS), supporting generation with arbitrary camera trajectories and timestamps, in natural scenes, conditioned on one or more images. With a novel architecture and sampling procedure, we enable training on a mixture of 3D (with camera pose), 4D (pose+time) and video (time but no pose) data, which greatly improves generalization to unseen images and camera pose trajectories over prior works that focus on limited domains (e.g., object centric). 4DiM is the first-ever NVS method with intuitive metric-scale camera pose control enabled by our novel calibration pipeline for structure-from-motion-posed data. Experiments demonstrate that 4DiM outperforms prior 3D NVS models both in terms of image fidelity and pose alignment, while also enabling the generation of scene dynamics. 4DiM provides a general framework for a variety of tasks including single-image-to-3D, two-image-to-video (interpolation and extrapolation), and pose-conditioned video-to-video translation, which we illustrate qualitatively on a variety of scenes. See https://4d-diffusion.github.io for video samples.
DJF is also affiliated with the University of Toronto and the Vector Institute.
@@
| Inputs | Extrapolate time |
| Inputs | Interpolate time |
@@ 1 Introduction
4D generative models cast novel view synthesis (NVS) as a form of image-conditioned multi-view generation with control of camera pose and time. They extend the impressive capabilities of generative image and video models, leveraging large-scale learning to enable image generation with consistent 3D scene geometry and dynamics. As such, they promise to unlock myriad generative tasks such as novel view synthesis from a single image, image to 4D generation, image(s) to video, and video to video translation. Nevertheless, designing and training 4D models that generalize well to natural scenes, with accurate control over camera pose and time, remains challenging.
While there has been interest in diffusion models for NVS (Watson et al., 2022; Liu et al., 2023b), much of this work has focused on isolated objects with cameras located on the viewing sphere around the object. A key barrier to generalization, both to arbitrary scenes and free-form camera poses, has been the lack of rich, posed training data. The predominant datasets, i.e., RealEstate10K (Zhou et al., 2018) and CO3D (Reizenstein et al., 2021), are restricted to narrow domains, with camera poses estimated from structure-from-motion (SfM) that are noisy and lack metric scale information. Consequently, state-of-the-art 3D NVS models with good generalization to date have focused on somewhat specific domains (e.g., CAT3D (Gao et al., 2024) for object-centric scenes, or PNVS (Yu et al., 2023a) for indoor scenes), and they exhibit issues with pose alignment, where generated images do not coincide with the specified target viewpoints. In the case of 4D NVS, said challenges around lack of large-scale posed training data and quality of pose estimation are exacerbated. To this end, we advocate co-training with unposed video, as diverse video is available at scale and contains valuable information about the spatiotemporal regularities of the visual world.
@@
To address these challenges, we introduce 4DiM, a diffusion model for NVS conditioned on one or more images of arbitrary scenes, camera pose, and time. 4DiM extends previous 3D NVS diffusion models in three key respects: 1) from objects to scenes; 2) to free-form camera poses specified in meaningful physical units; and 3) with simultaneous control of camera pose and time. 4DiM is trained on a mixture of data sources, including posed 3D images/video and unposed video, of both indoor and outdoor scenes. We make this possible through an architecture with FiLM layers (Dumoulin et al., 2018) that default to the identity function for missing conditioning signals (e.g., videos without pose, or static scenes without time), which we call Masked FiLM, and through multi-guidance, which leverages this kind of model to adjust guidance weights for different conditioning signals separately. We also introduce calibrated versions of datasets posed via COLMAP (Schönberger & Frahm, 2016); calibrated data makes it easier to learn metric regularities in the world, like the typical sizes of everyday objects and spatial relations, and it enables one to specify camera poses in meaningful physical units for precise camera control. As we later show, training on poses with unknown scale will (unsurprisingly) make scale itself stochastic when generating samples given one input image, whereas fixing this problem yields both improved fidelity and pose alignment.
To summarize our main contributions:
@@
We introduce 4DiM, a pixel-based diffusion model for novel view synthesis conditioned on one or more images of arbitrary scenes, camera pose, and time. 4DiM comprises a base model that generates 32 images at and a super resolution model that upsamples to ;
@@
We devise an effective data mixture for 4D models comprising posed and unposed video of indoor and outdoor scenes, yielding compelling results on camera control, time control, and promising zero-shot results controlling both simultaneously in video-to-video translation;
We propose Masked FiLM layers to enable training with incomplete conditioning signals, and multi-guidance which leverages such training to further improve sample quality at inference time;
@@
We create a scale-calibrated version of RealEstate10K to improve model fidelity and enable metric pose control;
@@
We include extensive evaluation and comparisons to prior work, including fixes to SfM-based metrics (He et al., 2024), and a novel keypoint distance metric to detect the presence of dynamics.
@@ 2 Related work
Subsequent work proposed attention mechanisms leveraging epipolar geometry to improve the 3D consistency of image-to-image diffusion models for NVS (Tseng et al., 2023), and more recently, finetuning text-to-video models with temporal attention layers or attention layers limited to overlapping regions (He et al., 2024; Wang et al., 2023a; Bahmani et al., 2024) to model the joint distribution of generated views and their camera extrinsics. These models however still suffer from several issues: they have difficulty with static scenes due to persistent dynamics from the underlying video models; they still suffer from 3D inconsistencies and low fidelity; and they exhibit poor generalization to out-of-distribution image inputs. Alternatives have been proposed to improve fidelity in multi-view diffusion models of 3D scenes, albeit sacrificing the ability to model free-form camera pose; e.g., see MVDiffusion and follow-ups (Tang et al., 2023; 2024) focusing on few, specific pose trajectories.
A concurrent body of related work on 3D extraction has recently emerged following DreamFusion (Poole et al., 2022), where instead of directly training diffusion models for NVS, new techniques for sampling views parametrized as a NeRF with volume rendering are proposed, such as Score Distillation Sampling (SDS) and Variational Score Distillation (Wang et al., 2024) (VSD). Here, a pre-existing diffusion model acts as a prior that drives the generation process. This enables, for example, text-to-3D using a text-to-image diffusion model. As demonstrated by ZeroNVS (Sargent et al., 2023), MVDream (Shi et al., 2023), ReconFusion (Wu et al., 2023) and CAT3D (Gao et al., 2024), using a pose-conditional diffusion model offers the advantage that the diffusion model can be conditioned on the sampled viewpoint for volume rendering during score distillation or NeRF postprocessing. Still, the extension of 3D extraction methods to 4D dynamic scenes and complex view-dependent reflectance phenomena remains challenging. In contrast, geometry-free diffusion models show encouraging results at scale with camera control, time control, and even their combination (Seitzer et al., 2023). Large-scale video models also provide encourage evidence of 3D geometrically consistent generation without an explicit rendering model (Brooks et al., 2024a; Li et al., 2024). Our success with 4DiM further supports this perspective, showing that one can simultaneously achieve fine-grained camera control and dynamics with a multiview diffusion model alone.
@@
4DiM uses a continuous-time diffusion model to learn a joint distribution over multiple views,
where are generated images, are conditioning images, are relative camera poses (extrinsics and intrinsics) and are scalar, relative timestamps.***4DiM does not require sequential temporal ordering as in video models as the architecture is permutation-equivariant over frames. All images (conditioning and generated) are processed by the diffusion model. Following Ho et al. (2020), we choose our loss function to be the error between the predicted and actual noise ( in their paper), though we use the L1 rather than L2 norm, because, like prior work (Saharia et al., 2022a; Saxena et al., 2023), we found that it improves sample quality. Our models use the “-parametrization” (Salimans & Ho, 2022), which helps stabilize training, and we adopt the noise schedules proposed by Kingma et al. (2021). Our current models process or 32 images at resolution ( is the combined number of conditioning plus generated frames). To this end, we decompose the task into two models similar to prior work (Ho et al., 2022a; b); we first generate images at , and then up-sample using a super-resolution model trained with noise conditioning augmentation (Saharia et al., 2022c). We also finetuned the 32-frame model to condition on 2 and 8 frames to enable more comparisons and example applications of 4DiM models.
@@
One particularly rich dataset for training 3D models is RealEstate10K (Zhou et al., 2018) (RE10K). It comprises 10,000 video segments of static scenes for which SfM (Schönberger & Frahm, 2016) has been used to infer per-frame camera pose, but only up to an unknown length scale. The lack of metric scale makes training more difficult because metric regularities of the world are lost, and it becomes more difficult for users to specify the target camera poses or camera motions in any intuitive units. This is especially problematic when conditioning on a single image, where scale itself otherwise becomes ambiguous. We therefore created a calibrated version of RealEstate10K by regressing the unknown metric scale from a monocular depth estimation model (Saxena et al., 2023). (For details, see Supplementary Material A.) The resulting dataset, cRE10K, has a major impact on model performance (see Section 4 below). We follow the same procedure to calibrate the LLFF dataset (Mildenhall et al., 2019) for evaluation purposes.
@@
Relatively little training data has both time and camera pose annotations. Most 3D data represent static scenes, while video data rarely include camera pose. Finding a way to effectively condition on both camera pose and time in a way that allows for incomplete training data is essential. We thus propose to chain Masked FiLM layers (Dumoulin et al., 2018) for (positional encodings of) diffusion noise levels, per-pixel ray origins and directions, and video timestamps. When any of these conditioning signals is missing (due to incomplete training data or random dropout for classifier-free guidance (Ho & Salimans, 2022)), the FiLM layers are designed to reduce to the identity function, rather than simply setting missing values to zero. This avoids the network from confusing a timestamp or ray coordinate of value zero with dropped or missing data. In practice, we replace the FiLM shift with zeros and scale with ones for masked signals. We illustrate the overall choices in Fig. 2. (For details see Supplementary Material C.)
@@
Steering 4DiM with the correct sampling hyperparameters is essential, especially the guidance weights for classifier-free guidance (Ho & Salimans, 2022) (CFG). In its usual formulation, CFG treats all conditioning variables as “one big variable”. In practice, placing a different weight on each variable is important; e.g., text requires high guidance weights, but high guidance weights on images can quickly lead to unwanted artifacts. We thus propose multi-guidance, where we generalize CFG to do exactly this, without making independence assumptions between conditioning variables. In essence, multi-guidance generalizes 2-guidance originally proposed in InstructPix2Pix (Brooks et al., 2023) to variables. We start from a classifier-guided formulation (Dhariwal & Nichol, 2021), where we wish to sample with conditioning signals, , using the score
@@
If we train our model such that it drops out only , or and , … , or all of , we can sample with different guidance weights on each conditioning variable . 4DiM is trained to drop out conditioning signals with probability 0.1, and it drops either , and , or all of , and . This is a natural choice, since poses and timestamps without corresponding conditioning images do not convey useful information. For best results, 4DiM uses a guidance weight 1.25 on conditioning images and weights of 2.0 on camera poses and timestamps. We include an ablation study in our Supplementary Material B that demonstrates how multi-guidance provides additional knobs to push the Pareto frontiers between fidelity and dynamics / pose alingnment.
@@ 4 Evaluation
Evaluating 4D generative models is challenging. Typically, methods for NVS are evaluated using generation quality. In addition, for pose-conditioned generation it is desirable to find metrics that capture 3D consistency (rendered views should be grounded in a consistent 3D scene) and pose alignment (the camera should move as expected). Furthermore, evaluating temporal conditioning requires some measure that captures the motion of dynamic content. We leverage several existing metrics and propose new ones to cover all these aspects, as described below.
@@
FID (Heusel et al., 2017) is a key metric for image quality, but if used in isolation it can be uninformative and a poor objective for hyper-parameter selection. For example, Saharia et al. (2022b) found that FID scores for text-to-image models are optimal with low CFG weights (Ho & Salimans, 2022), but text-image alignment suffers under this setting. In what follows we report FID as well as the improved Frechet distance with DINOv2 features (FDD) (Oquab et al., 2023), which appears to correlate with human judgements better than InceptionV3 features (Szegedy et al., 2016). For video quality, we also report FVD scores (Unterthiner et al., 2018).
@@
The work of Yu et al. (2023a) introduced the thresholded symmetric epipolar distance (TSED) to evaluate 3D consistency, as a more computationally efficient alternative to training separate NeRF models (Mildenhall et al., 2021) on every sample (e.g., like (Watson et al., 2022)). First, SIFT is used to obtain keypoints between a pair of views. For each keypoint in the first image, we can compute the epipolar line in the second image and its minimum distance to the corresponding keypoint. This is repeated for each keypoint in the second image to obtain a SED score. A hyperparameter threshold is selected, and the percentage of image pairs whose median SED is below the threshold is reported. Below, we report TSED with a threshold of 2.0 and include the TSED score of ground truth data, as poses in the data are noisy and do not achieve a perfect score.
Pose alignment (revised metric: SfM distances)
We follow He et al. (2024) to quantify pose alignment, i.e., by running COLMAP pose estimation on the generated views, and comparing the camera extrinsics predicted by COLMAP to the target poses. To ensure the results are reasonable metrics, however, we modify the procedure of He et al. (2024) in several ways. First, due to the inherent scale and rotational ambiguity of COLMAP, we observe that it is necessary to align the estimated poses to the original ones before comparing their differences as described above. We also report the relative error in camera positions (SfMDpos) and the angular deviation of camera rotations in radians (SfMDrot) to account for scale variation across scenes. Finally, to get best results from COLMAP, we also feed it the input views. Please see our Supplementary Material E for more detail. We refer to this set of metrics as SfM distances hereon. Like we do for TSED scores, we also report the SfM distances for the ground truth data.
@@
One common failure mode that we observed early in our work is a tendency for models to copy input frames instead of generating temporal dynamics. Prior work has also noted that good FVD scores can be achieved with static videos (Ge et al., 2024). We therefore propose a new metric called keypoint distance (KD), where we compute the average motion of SIFT keypoints(Lowe, 2004) across image pairs with matches. We report results on generated and reference views to assess whether generated images have a similar motion distribution.